a steel plug has a diameter of 5cm at 30 degrees celsius. At what temperature will it fit exactly into a hole of constant diameter 4.997cm?
1
Expert's answer
2020-06-30T18:24:48-0400

By the definition of the linear thermal expansion we have:


"\Delta L = \alpha L_0 \Delta T,"

here, "\Delta L = L - L_0" is the increace in diameter of the steel plug when it is heated, "L_0 = 0.04997 \ m" is the initial diameter of the steel plug at some temperature "T_0", "L = 0.05 \ m" is the diameter of the steel plug at "T = 30 \ ^{\circ}C", "\alpha = 11 \cdot 10^{-6} \ ^{\circ}C^{-1}" is the linear expansion coefficient for steel, "\Delta T = T - T_0" is the change in temperature.

From this formula we can find the change in temperature:


"\Delta T = \dfrac{L - L_0}{\alpha L_0},""\Delta T = \dfrac{0.05 \ m - 0.04997 \ m}{11 \cdot 10^{-6} \ ^{\circ}C^{-1} \cdot 0.04997 \ m} = 54.6 \ ^{\circ}C."

Finally, we can calculate the temperature at which the steel plug will fit exactly into a hole of constant diameter:


"T_0 = T - \Delta T = 30 \ ^{\circ}C - 54.6 \ ^{\circ}C = -24.6 \ ^{\circ}C."

Answer:

"T_0 = -24.6 \ ^{\circ}C."